# Compressed Feature Quality Assessment: Dataset and Baselines

### v1

@@ 5.1. Baseline CFQA Metrics
To cover diverse types of distortions and similarity relationships between features, we select three representative signal-based metrics: MSE, Cosine Similarity, and Centered Kernel Alignment (CKA)  (Kornblith et al., 2019).
@@
  • Cosine Similarity: Cosine similarity measures the angular difference between feature vectors, normalized by their magnitudes. This metric captures directional alignment in the feature space, which is often more robust to scale distortions and is considered to better reflect semantic similarity in high-dimensional embeddings (Radford et al., 2021).
  • CKA: CKA measures similarity between two features using normalized HSIC (Hilbert-Schmidt Independence Criterion). Unlike MSE or cosine similarity, CKA captures structural relationships and global correlation patterns between two features, making it suitable for comparing multi-token or multi-patch feature distributions as commonly found in vision transformers.
These metrics form diverse baselines from element-level to structure-level distortion metrics.
@@
Figure 2. PLCC distribution visualization of MSE, cosine similarity, and CKA. The first, second, and third rows correspond to the Cls, Seg, and Dpt tasks, respectively. The first, second, third, and fourth column correspond to the HM, VTM, multi-task-trained Hyperprior, and task-specific-trained Hyperprior codecs.
@@ 5.2. Evaluation Protocol
We define a evaluation protocol to assess how well a CFQA metric predicts the semantic distortion of compressed features. PLCC and SROCC are used in our experiments.
@@
Specifically, for each original feature $\mathbf{f}_{i}$, we compute these two metrics on its 10 predicted quality scores and 10 true semantic distortion. Each metric is evaluated on all three tasks and across all codecs.
@@ 5.3. Rate-Accuracy Performance Analysis
Table 2 presents the rate-accuracy performance of the handcrafted codecs. For all three tasks, both codecs exhibit a broad range of bitrates and corresponding accuracy levels, spanning from nearly lossless reconstruction to significant semantic degradation. This wide span underscores their utility in effectively simulating traditional coding distortions. However, the Dpt task shows a narrower variation in both bitrate and performance. This can be attributed to the use of multi-scale features, which include lower-level layers with higher redundancy and reduced semantic abstraction, resulting in less fluctuation in both bitrate and distortion.
Table 3 shows the rate-accuracy performance of two learning-based codecs. In contrast to handcrafted codecs, the learning-based codecs exhibit more varied performance behaviors across tasks. Specifically, the multi-task codec outperforms the task-specific codec in Cls, whereas the task-specific codec demonstrates superior performance in Dpt. This suggests that incorporating fine-grained task features enhances the compression training of coarse-grained features, while fine-grained features are more effectively trained using a single distribution tailored to the task.
@@ 5.4.1. Average Performance Analysis
Among the three baseline metrics, cosine similarity consistently demonstrates a higher degree of linearity and monotonicity in relation to true semantic distortion. This is due to the metric’s ability to capture angular relationships between feature vectors, making it particularly effective for high-dimensional feature sets such as those produced by ViT.
@@
Among the three tasks, Seg presents the most complex and difficult-to-fit distortions. For most codecs, the baseline metrics show poorer fitting for Seg compared to Cls and Dpt. This reflects the higher semantic complexity in Seg, which is more challenging for these metrics to capture accurately.
@@ 5.4.2. Distribution Analysis
Among the three baseline metrics, cosine similarity exhibits the most concentrated PLCC distributions. This observation aligns with its superior average PLCC values, as reported in Table 2 and Table 3, indicating that cosine similarity provides a more stable quality estimation across varying conditions.
@@
5.5. Limitations
While the evaluated metrics demonstrate moderate performance in certain tasks and codecs, they also exhibit significant limitations. First, signal-based metrics like MSE lack task sensitivity and often fail to reflect meaningful semantic degradation, especially in segmentation and generation tasks. Second, cosine and CKA metrics, although capturing structural alignment, are sometimes unstable under nonlinear distortions introduced by learned codecs. Finally, none of the metrics generalize consistently across all tasks and compression strategies, suggesting that handcrafted similarity alone may not suffice for universal FQA. These limitations highlight the need for more adaptive, task-aware, or learned semantic quality estimators, which we consider for future work.
@@ 6. Conclusion
This paper introduces the concept of Compressed Feature Quality Assessment (CFQA), a crucial area for evaluating the semantic degradation of compressed features in systems where features are transmitted, stored, and reused. We present the first benchmark dataset for CFQA, which lays the groundwork for further research in this field. We assess the widely used metric in CFQA and provide insights for interested researchers. Future research will focus on developing adaptive CFQA metrics that generalize across tasks and feature coding strategies.

### v2

@@ 5.1. Baseline CFQA Metrics
To cover diverse types of distortions and similarity relationships between features, we select three complementary metrics: MSE, cosine similarity, and CKA  (Kornblith et al., 2019). These metrics collectively span from local element-wise to global structural comparisons.
@@
  • Cosine similarity: Cosine similarity measures the angular difference between feature vectors, normalized by their magnitudes. This metric captures directional alignment in the feature space, which is often more robust to scale distortions and is considered to better reflect semantic similarity in high-dimensional representations (Radford et al., 2021).
  • CKA: CKA measures similarity between two features using normalized HSIC (Hilbert-Schmidt Independence Criterion). It captures higher-order statistical relationships, making it particularly suitable for comparing architectural differences and global feature patterns.
@@
Figure 2. PLCC distribution visualization of MSE, cosine similarity, and CKA. The first, second, and third rows correspond to the Cls, Seg, and Dpt tasks, respectively. The first, second, third, and fourth columns correspond to the HM, VTM, multi-task-trained Hyperprior, and task-specific-trained Hyperprior codecs.
@@ 5.2. Evaluation Protocol
We define an evaluation protocol to assess how well a CFQA metric predicts the semantic distortion of compressed features. PLCC and SROCC are used in our experiments.
@@
Specifically, for each original feature $\mathbf{f}_{i}$, we compute these two metrics on its 10 predicted quality scores and 10 true semantic distortions. Each metric is evaluated on all three tasks and across all codecs.
@@ 5.3. Rate-Accuracy Performance Analysis
Table 2 presents the rate-accuracy performance of the handcrafted codecs. For all three tasks, both codecs exhibit a broad range of bitrates and corresponding accuracy levels, spanning from nearly lossless reconstruction to significant performance degradation. This wide variation highlights their effectiveness in simulating traditional coding distortions. However, the Dpt task exhibits less variation in both bitrate and performance. This difference stems from the use of multi-scale features, which include lower-level layers with higher redundancy and reduced semantic abstraction. These characteristics lead to smaller fluctuations in both bitrate and distortion.
Table 3 shows the rate-accuracy performance of two learning-based codecs. Unlike the handcrafted codecs, the learning-based codecs show more diverse performance patterns across tasks. The multi-task Hyperprior codec achieves better performance than the task-specific Hyperprior codec for Cls, while the task-specific Hyperprior codec performs better in Dpt. These results indicate that incorporating fine-grained task features improves the codec optimization of coarse-grained features, whereas fine-grained features benefit more from training on a single task-specific distribution.
@@ 5.4.1. Average Performance Analysis
Among the three baseline metrics, cosine similarity consistently demonstrates a higher degree of linearity and monotonicity in relation to ground-truth semantic distortion. This is due to the metric’s ability to capture angular relationships between feature vectors, making it particularly effective for high-dimensional features such as those produced by ViTs.
@@
Among the three tasks, Seg presents the most complex and difficult-to-fit distortions. For most codecs, the baseline metrics show poorer fitting for Seg compared to Cls and Dpt. This reflects the higher semantic complexity in Seg, which is more challenging for these metrics to capture.
@@ 5.4.2. Distribution Analysis
Among the three baseline metrics, cosine similarity exhibits the most concentrated PLCC distributions. This observation aligns with its superior average PLCC values, as reported in Table 4, indicating that cosine similarity provides a more stable semantic distortion measurement across varying conditions.
@@
5.5. Discussion
Our evaluation reveals three key limitations of current CFQA metrics. First, signal-based metrics like MSE show limited task sensitivity and often correlate poorly with actual semantic degradation. Second, while cosine similarity and CKA capture structural relationships, they demonstrate instability when handling nonlinear distortions from learned codecs. Most importantly, none of these metrics achieves consistent performance across all tasks and compression methods, indicating that conventional similarity measures alone cannot provide universal CFQA solutions. These findings underscore the necessity for developing more adaptive, task-aware quality estimators – a direction we plan to explore in future work.
@@ 6. Conclusion
This paper introduces the concept of Compressed Feature Quality Assessment, a crucial research area for evaluating the semantic distortion of compressed features in systems where features are transmitted, stored, and reused. We present the first benchmark dataset for CFQA, which lays the groundwork for further research in this field. We assess the widely used metrics in CFQA and provide insights for interested researchers. Moving forward, we will focus on developing adaptive CFQA metrics capable of generalizing across diverse tasks and coding strategies.
Acknowledgements.
This work was supported by the Ministry of Education of Singapore under Grant T2EP20123-0006. We acknowledge the support of GPU cluster built by MCC Lab of Information Science and Technology Institution, USTC.